how many moles of (CH3)3NH+ ARE IN 6.0 g of (CH3)3NH+?
1
Expert's answer
2013-01-29T10:14:09-0500
To find the amount of the substance (CH3)3NH+ to calculate its molar mass:
M((CH3)3NH+) = (12+3)*3 + 14+1 = 60 g/mol
n((CH3)3NH+) = m/M
m((CH3)3NH+) = 6g
n((CH3)3NH+) = 6g/60 g/mol = 0.1 mol

Answer: n((CH3)3NH+) = 0.1 mol